OpenStudy (shelbywyatt):

is y/0=n/0 n=any number with a absolute value more or equal to 0 y= any number NOT "n", with a absolute value more or equal to 0 Is this true?

OpenStudy (zzr0ck3r):

no anything/0 is not defined and so the statement y/0= is not even a real thing.

OpenStudy (skullpatrol):

This is NOT true because the expression a/0 has no meaning in the set of real numbers. $$\Huge \dfrac{a}{0}=a∙\dfrac{1}{0}$$ But 0 has no reciprocal because 0 times any number is 0, not 1 as required by a • (1/a) = 1. Do you see the contradiction? :-)

OpenStudy (shelbywyatt):

yeah :)

OpenStudy (zzr0ck3r):

Is there any field where you can divide by the additive identity?

OpenStudy (shelbywyatt):

no? I think